Paolo's monthly salary is 1 3/7 that of Francesca's. Francesca's monthly salary is 7/5 of Dante's. What fraction of Paulo's earnings is Dante's?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

Paolo's monthly salary is 1 3/7 that of Francesca's. Francesca's monthly salary is 7/5 of Dante's. What fraction of Paulo's earnings is Dante's?
.
Let D = Dante's salary
and P = Paolo's monthly salary
then
(7/5)D = Francesca's salary
.
So, Paolo's monthly salary is:
(1+3/7)(7/5)D =P
(7/7+3/7)(7/5)D =P
(10/7)(7/5)D =P
(10/1)(1/5)D =P
(2/1)(1/1)D =P
(2)D =P
D = (1/2)P
.
So, Dante's salary is 1/2 of Paulo's

Answer by stanbon(75887)

Paolo's monthly salary is 1 3/7 that of Francesca's. Francesca's monthly salary is 7/5 of Dante's. What fraction of Paulo's earnings is Dante's?
-----------------------
---
P = (10/7)F
F = (7/5)D
-----------------
Substitute for "F" in the 1st equation to get:
P = (10/7)(7/5)D
P = 2D
D = (1/2)P
----
Dante's earnings is 50% of Paulo's.
=====================================
Cheers,
Stan H.